Express 54069,269 in words; place value; and in terms of the powers of 10.
1
Expert's answer
2020-05-05T20:15:09-0400

54069269 in words is "fifty-four million, sixty-nine thousand, two hundred, sixty-nine.

54069269 in place value:-

"\begin{matrix}
 Ten-Million & Million & Hundred-thousands & Ten-thousands \\
 5 & 4 & 0 & 6 
\end{matrix}" "\begin{matrix}
 Thousand & Hundreds & Tens & Ones\\
 9 & 2 & 6 & 9
\end{matrix}" .

54069269 in power of 10 = "5.4069269 \times 10^7".